solve using the principles together 8y-7=23-2y...Solution is Y=
Answer by EricL(29)

Hello!
Problem: 8y-7=23-2y
Move all terms that don't contain y to one side of the equation.
10y-7=23
10y=30
y=3
--------------------
If you have any questions about this problem, feel free to ask :)
--------------------
- Eric L